x^2+2x-11=0 ??
solve each equation by completing the square thanks for the help :)))

Answer by lwsshak3(11628)

x^2+2x-11=0 ??
solve each equation by completing the square
(x^2+2x+1)-11-1=0
(x+1)^2-12=0
(x+1)^2=12
take sqrt of both sides
x+1=±√12
x=-1±√12=-1-√12 or -1+√12